Good Saturday morning Tutors.
I need help to write the slope-intercept equation for the line that passes through (-5,10) and (-4,3).

Answer by nerdybill(7384)

write the slope-intercept equation for the line that passes through (-5,10) and (-4,3)
.
Begin by finding the slope between (-5,10) and (-4,3)
m = (y2-y1)/(x2-x1)
m = (3-10)/(-4-(-5))
m = (-7)/(-4+5)
m = (-7)/(1)
m = -7
.
Use the slope (-7) and either one of the provided points (-4,3), plug into the "point-slope" form:
y - y1 = m(x - x1)
y - 3 = -7(x - (-4))
y - 3 = -7(x + 4)
y - 3 = -7x - 28
y = -7x - 25 (this is what they're looking for)